OpenStudy (anonymous):

What is the mathematical equation of twice a number is 6?

OpenStudy (gorv):

let the number is x 2*x=6

OpenStudy (anonymous):

do you have another answer?

OpenStudy (phi):

twice a number is 6 twice a number means the same thing as two times a number is 6 change two to 2 2 times a number is 6 change "a number" to a *variable name* People use "x" for the "an unknown number" 2 times x is 6 change "times" to a multiply sign 2 * x is 6 "is" almost always changes to an equal sign = 2 * x = 6 which is the equation you want. In algebra, we don't bother to write the multiply sign, so we could also write 2x = 6

OpenStudy (anonymous):

ok thankyou

OpenStudy (anonymous):

is it right that twenty-five decreased by twice a number is twelve the answer is 25-2x=12? or it is wrong?

OpenStudy (phi):

twenty-five decreased by twice a number is twelve to figure out what "decreased by" means I try a simple problem 2 decreased by 1 gives 1 or 3 decreased by 1 gives 2 so it must mean subtraction. now to interpret: twenty-five decreased by twice a number is twelve 25 - 2 * x = 12 or 25 -2x=12 yes you have the correct answer.

OpenStudy (anonymous):

ok thankyou !